CTRC - The California Trolley and Railroad Corporation.
Preserving the historical heritage of railroading and trolleys in the Santa Clara Valley of California.

Lost Postings of 2013

Several items fell through the cracks:

June 29, 2013

100 degree temps only slowed us down a little!

Machining of the bushing for the left rear side rod number 3 driver was completed during the week.  This morning it was pressed into place.  Measurements were then taken to determine if the buffering ring was of the proper width.  Unfortunately the bushing was to long.

The buffering ring prevents the bushing from being pressed out of position in the rod/bushing come in contact with the washer which is between the rod and nut.  If the bushing moves it will wear inside the rod and eventually become sloppy in the rod.  The buffering ring is pressed on the end of the bushing.  If fit is proper, the end of the bushing should be slightly lower inside the ring.

With the bushing being to long, it was pressed out and trimmed.  Bushing was then pressed back into the rod.  Buffering ring pressed into place.  Rod was then moved to the locomotive.

There were no issues getting number 3 pin to accept the rod.  Problems came about with alignment of the front end of the rear rod to the knuckle at the rear of the front side rod.  Number 3 driver was needed to come forward about one eighty of inch.

A 20 ton bottle jack was set between the locomotive frame and a large spoke.  Heavy gear oil was brushed between the rail and wheel.  Jack was then used to rotate the wheel a very small amount.  Surprisingly the wheel did rotate rather easily considering the weight the wheel bears!

There was a bit of "house" keeping performed.  Securing of the galvanized steel sheets that were hung on the front shop doors was completed.  This coming week a coat of primer may be applied.  The back wall of the outside shop will now be covered with the remaining galvanized material. Presently a large plastic tarp covers this area.  This tarp is about to fail and would be expensive to replace.

July 27, 2013

Movement forward on several fronts.

The bushing for the right main rod is coming along.  The inside and outside diameters are complete.  Bushing now must be cut on band saw, and finished.  Locating and drill of the many grease holes will follow.

Cleaning of the main rods is complete.  The last side was finished today.

One set of caboose steps was sanded and coat of paint applied.

During this past week, a good deal of time was allocated to the locomotive's steam dome cover.  The present cover is to thin.  It was noted on RYPN that the TVRR had purchased a new cover for their 4501.  Contact was made with TVRR.  They used the vendor that was being considered to form 2479's new cover.  Cost was in the vicinity of $500.  This is good news.  A quote request for a new cover for 2479 will be submitted this coming week.

August 10, 2013

Machining of the last two bushings continues.  21 grease holes have been drilled and chamfered in the bushing for the right main.  Chamfering of the holes on the inner surface and cutting a radius on the inside edge remain.

Machining of the left main bushing is about 50% complete.  The inner and outer diameters are complete.  The end of the future bushing that was clamped in the jaws of the Le Blond chuck was cut on a band saw.  A bunch of work remains.

Repairs to the cab are ongoing.  Welding in a section of the engineer's rain gutter continues. Some good head way was made today.

Hyster forklift has had problems with the right front tire loosing pressure.  Inner tube was patched twice but leak continues.  New inner tube was installed today.  This little forklift is handy when heavy items have to be moved into the outside shop.

Improvements to the outside shop continue.

August 17, 2013

This morning machining of the right main rod bushing was completed and the main rod installed.  Installation went rather smooth.  The rod will have to be removed once the pistons are installed to check limits of piston travel.  It was felt that bushing and rod surfaces are better protected by storing them on the locomotive.

Bushing for L2 main rod driver pin is about 60% complete.  Locating, drilling and chamfering of grease holes is next.

Art has machined two new grease fittings for the driver pins.  At the moment CTRC does not own an Alemite stick grease gun.  It also appears that there are no vendors for the large button head fittings previously used by SP.  Art has machined an adaptor with a 1/2" button head grease fitting.  We do have an pneumatic grease gun capable of pushing grease threw the pin and bushing passages.  Two of these adapters were installed and tested.  Looks like rods and pins can now be lubricated during the remainder of the restoration.  Art is going to machine adapters for the remaining four pins.

Problems today with our portable battery charger.  Some tinkering and unit was again serviceable.

We are studying a different way of attempting to roll the locomotive.  There is concern that the present method of forklifts and chains is going to cause serious damage to one or both forklifts.

During the week material for a new steam dome cover was placed on order.  Total cost for this cover is $417.00.  Cover will be near identical to the present cover but heavier in all of the necessary places.  Object should arrive late September.

August 24, 2013

Machining of the last bushing, L2 main rod, is complete! Today we located, drilled and chamfered the 21 grease holes.  With this work out of the way, the bushing was again test fit.  Next Saturday morning the left main rod will be mounted.  For those interested in observing, work should being about 9:30 AM.

Another big "Thank You!" to Karen and Jerry Lane for the grant that allowed CTRC to move a big step forward in the restoration of X-SP-2479.

Adaptors for 1/2 button head grease fittings have been machined.  This will allow grease to be injected into the pins and rods during the remaining restoration period.  Keeping grease in the rods will prevent corrosion from forming in the bottom of the rods.

August 31, 2013

Today the left main rod was returned to the locomotive.  Installation of the rod and main pin bushing went rather well.  Several hours were spent cleaning and repairing the threads in the crosshead pin for the grease adapter.

In the above image the cross head pin is sitting on a tie adjacent to the rear wheel of the engine truck.

Repairs to the rain gutter on the engineer's side of the cab are complete.  A section of the gutter was cutter away and a new section welded into place.

September 7, 2013

Day was spent forcing grease into a couple of rod pins.  Finally were successful getting grease into both cross head/main rods and the main rod bushing on left side.  I'm not sure about grease making its way back to the main pin's side rod bushing on the left side.

Believe most of the work on the engineer's side rain gutter is complete.

September 28, 2013

Small crew today.

We were unable to verify that grease was flowing threw the number 2 left driver pin to the bushing for the front rod.  It was deemed necessary to pull the left main and front side rod.  The grease passages in the left main pin were cleaned and grease flow was verified.  The rods were then installed.  Process only required 3.5 hours.

The new steam dome cover is suppose to ship Monday from Fort Worth, TX.  I would expect it to arrive late in the week.

October 1, 2013

New cover arrived this morning.  Initial measurements are excellent.

October 5, 2013

Several significant tasks under way.

Several of the machining steps on the new steam dome cover have been completed.  The gasket side of the cover was machined flat earlier in the week.  Today the top side of the flange was machined flat and a center hole was drilled and tapped.  The preceding steps were completed on the barn's Leblond lathe.  (We could not have completed a good number of tasks without that machine!) The cover will now be moved to the restoration site, where 24 one inch holes will be drilled.

Our welder has returned from an extended "vacation" in Pennsylvania.  Welding upon the firebox wrapper sheet on the fireman's side has again begun.  A good number of pounds of welding rod is going to be required in building up thin areas.

More of the cab sides have been wire brushed and primer applied.

October 19, 2013

Small but productive crew.

Firebox wrapper sheet pad welding- Light grinding and sanding was performed on the welded area on the fireman's side of the firebox wrapper sheet.  Ultra sonic measurements are much improved but a second pass is going to be required.  That process started today.  Unfortunately the welder ran out of the needed rod with about 1/3 of the work complete.

Steam dome cover- After 8 attempts and adjustment in hole location, the new steam dome cover finally accept the dome's studs.  It appears that at least 12 studs will have to be replaced to provide proper length.

Cab- The cab's side vents were removed and hinges cleaned.  Vents now move with ease.

Winterizing locomotive- Now that rod bushings are in place, new plastic tarps were acquired and set in place to provide protection from up coming seasonal rains.

Thanks to all that participated.

November 2, 2013

As in the past, there is progress.  Over the past two weeks, 22 of the 24 steam dome studs were extracted from the dome.  This was necessitated due to the thickness of the new cover.  Studs were removed by welding a nut on to the stud.  An large air impact wrench was then used to unscrew the stud.

Machining of new studs is well underway.  Three of the new studs are already in the dome.  If all goes well, all studs should be back in the dome by close of business next Saturday.

Pad welding on a large area of the firebox continues.  The area that is being built up is about 12 inches by 5 inches.  Two passes have been completed.  The area requires grinding and ultra sonic after which ultra sonic measurements can be taken.  A third weld pass my be required in a smaller area.

Cleanup upon the exterior of the cab continues.  If weather holds, a primer coat will be applied to the exterior next Saturday.

November 9, 2013

Several members that have not been available fir some time joined today's work crew and helped make this a extremely productive session.

Today the new steam dome cover was set into place.  This is a significant mile stone.  In September of 2005 (yes, eight years one month past) calculations based upon ultra sonic measurements revealed the old cover to thin.  Regardless of how the calculations were manipulated, the results were always the same, cover not serviceable at 210 PSI.

Machining on cover was completed several weeks ago, but the thicker cover required replacement of 22 of 24 dome studs.  Material for new studs was on hand.  Longer studs were machined.  Today those studs were set tight.  The new cover slipped into place after a couple of small "adjustments" were performed.

Every 3 or 4 months the locomotive and tender are rolled to keep journals lubricated.  In the past, two forklifts were employed.  The Clark forklift although large is not suitable for this task.  Its not geared low enough.  Over the years of CTRC service, several components in the drive line have failed.  Continued use this machine in this manner eventually lead to another costly if not catastrophic failure.

Today we experimented using a jeep winch and pulleys to pull the locomotive.  The Wiggins lift was required to assist climbing a rail joint other wise the move went well.  There are a couple of issues to resolve.  It now appears that the only use of the Clark during future moves of the locomotive is as an anchor for the jeep.

The stairs that are used to access the locomotives cab area received some much needed help.  A new brace was added and preparations were made to apply a new coat of paint.

The surfaces of the turntable and stand for the water tank were cleared of leaves from the pepper tree.  Request has been made to fairgrounds staff to have the tree pruned which may happen in the next week or two.

November 26, 2013

The OSH Box car that sat along Azuerais Ave.  for 50 plus years was moved today to San Jose's History Park.

In the coming weeks the display track that SP1215 and MoPac caboose sit upon will be lengthened and the boxcar set between the locomotive and caboose.

The move would not have gone as nice as it did without the help of Tom Anderson and Larry Murchison.  On Saturday, Tom helped preparing the car and getting cribbing materials pre-positioned.  Today, Larry helped positioning crib material to set the car body upon.

November 30, 2013

Measurements were taken of the inside diameter of the cylinder that holds the left front valve cage. This was needed to update the valve cage drawing.  Once an accurate drawing has been created, the drawing can be submitted to vendors for machining of a new cage.

Performed NDT on one of three of the locomotive's air reservoirs.  News is not good.  A deep gouge and waste area was found on visual inspection.  The area was polished to form a flat spot for the transceiver of the Ultra Sonic Thickness gauge.  Measured thickness was .177.  Calculations call out a minimum thickness of .223 for a 110 PSI air system.  By Part 230 rule, air reservoirs can not be repaired.  This tank is not fit for service.

Over the coming weeks the remaining tanks will be inspected.

December 7, 2013

Not much accomplished today.  Lower temperature and moderate wind kept all hands close to heaters.

Some time was spent examining the second of three air tanks.  Again news is not good. A waste area of substantial depth about size of a dime was located on the largest tank, 90" length.  Two tanks now not serviceable.

Drawings will be constructed.  Once reviewed, drawings will be sent off to ASME shops to obtain quotes.

December 15, 2013

A bit of progress yesterday.

Cab - Furring strips along the front and rear of the cab roof that were cut out and installed years ago by the trolley group were installed with more appropriate hardware.  Repairs to the last segment of damaged fireman's rain gutter was started.

Cab will soon be ready for wood liner.

Left Front Valve Cage - Very accurate measurements were taken of the inside surface in the left front valve's cylinder.  These numbers are needed to complete a drawing for the new valve cage.  Several other small issues remain with this drawing.  Drawing should be ready for quotes in early January.

Air Reservoirs - Third air tank was removed from the locomotive to enable inspection.  A generic drawing has been started to be used to obtain quotes.  Hopefully all work will be complete and ready for quotes by the end of January.

Boiler - Grinding and polishing of an area on the fireman's side of the fire box that has seen several layers of pad welding applied.  This area is about 5 inches by 24 inches.  Ultra sonic (UT) measurements were taken.  Welded area now meets required minimum thickness.  There is still a significant area to be welded along top and one side.  Fortunately there are no restrictions on the size of an area that can be pad welded in a stayed area, otherwise this would have been a very expensive repair.  There are several other areas requiring pad welding.  These areas are smaller then the section presently under repair.

A UT measurement was taken in each square in the above image and is recorded.  A bit more grinding and polishing is needed to complete the work.


Sponsored Links